OpenStudy (anonymous):

The coordinates of the vertices of two rectangles, Rectangle ABCD and Rectangle PQRS are given below. Rectangle ABCD – A(0, 8), B(8, 8), C(8, -4), D(0, -4) Rectangle PQRS – P(0, 4), Q(4, 4), R(4, -2), S(0, -2) Maria scaled both the rectangles about their centers to create two congruent rectangles A′B′C′D′ and P′Q′R′S′. By which factor did she most likely scale Rectangle ABCD and Rectangle PQRS?

OpenStudy (anonymous):

Rectangle ABCD by and Rectangle PQRS by 2 to create congruent rectangles of dimensions 4 x 3 units Rectangle ABCD by 2 and Rectangle PQRS by 3 to create congruent rectangle of dimensions 3 x 2 units Rectangle ABCD by and Rectangle PQRS by 4 to create congruent rectangles of dimensions 4 x 3 units Rectangle ABCD by and Rectangle PQRS by to create congruent rectangles of dimensions 2 x 3 units

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the distance from A to B?

OpenStudy (anonymous):

8 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the distance from P to Q

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so you want to multiply each distance by some number to get them to the same number

OpenStudy (anonymous):

i guess? I don't really understand this

jimthompson5910 (jim_thompson5910):

let's say for instance the answer is B (it's not, but let's say it is) Rectangle ABCD by 2 and Rectangle PQRS by 3 to create congruent rectangle of dimensions 3 x 2 units so we would multiply the distance of 8 by 2 to get 16 then multiply the distance of 4 by 3 to get 12 but 16 doesn't equal 12, so that's why B isn't the answer

OpenStudy (anonymous):

ohh okay so let me try A Rectangle ABCD by 1/2 and Rectangle PQRS by 2 to create congruent rectangles of dimensions 4 x 3 units

OpenStudy (anonymous):

8*1/2=4 4*2=8 so it's not A/

jimthompson5910 (jim_thompson5910):

yep 8*(1/2) = 4 4*2 = 8 doesn't work (4 = 8 is false)

OpenStudy (anonymous):

C Rectangle ABCD by 1/4 and Rectangle PQRS by 4 to create congruent rectangles of dimensions 4 x 3 units

OpenStudy (anonymous):

8*1/4=2 4*4=16

OpenStudy (anonymous):

no right?

jimthompson5910 (jim_thompson5910):

correct, that doesn't work either

OpenStudy (anonymous):

D Rectangle ABCD by 1/4 and Rectangle PQRS by 1/2 to create congruent rectangles of dimensions 2 x 3 units

OpenStudy (anonymous):

is it D?

jimthompson5910 (jim_thompson5910):

8*(1/4) = 2 4*(1/2) = 2 works (2 = 2 is true)

jimthompson5910 (jim_thompson5910):

so because can show it's true and we've eliminated everything else, it has to be D

OpenStudy (anonymous):

yay! thank you :D

jimthompson5910 (jim_thompson5910):

yw